OpenStudy (anonymous):

Need help with some questions on finding the probability of different outcomes. WILL FAN AND MEDAL.

OpenStudy (anonymous):

hit me wit dem

OpenStudy (anonymous):

I think 10/12 are there answer choices will explain

OpenStudy (anonymous):

5/32

OpenStudy (anonymous):

nop im wrong sorry

OpenStudy (anonymous):

is 10 /12

OpenStudy (welshfella):

total possible outcomes = 6*6 = 36 a 1 on each dice will give a sum of 2

OpenStudy (welshfella):

P( rolling 2 1's0 = 1/36

OpenStudy (welshfella):

P (rolling 2 1's) = 1/36

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

so is it 10/12 or not? o.o

OpenStudy (welshfella):

no

OpenStudy (anonymous):

So 35/36?

OpenStudy (welshfella):

yes there is only one way to get a sum of 2 so there are 36-1 ways of NOT getting a sum of 2

OpenStudy (welshfella):

ok?